If the compound amounts of a sum of money in 2 yrs and 3 yrs are Rs 13000 and Rs 14300 respectively, find the rate of interest.

Grade:10

1 Answers

Vikas TU
3 years ago
Difference between c. i. of any two consecutive years is equal to si calculated on small ci for year

14300-13000=P.r.t/100
1300=13000*r*1/100
1300*100/13000=r
10 % =r